# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 1.3. Overview and main technical components
The careful treatment of harmonic bundles on the regular locus presented above is necessary in the context of compact Kähler klt spaces because Mochizuki’s theory of tame harmonic bundles on quasi-projective manifolds [83, 84, 85]—which crucially underlies the construction of $\mu_{X_{\mathrm{reg}}}$ in the projective case [39]—has not yet been established for general quasi-Kähler manifolds.
It should be noted that the admissibility of the limiting Hermitian–Einstein metric constructed in Section 3 remains unknown (cf. Question 3.6). The key point of Steps 2 and 3 is that klt singularities are sufficiently mild to allow the characterization of polystable Higgs bundles satisfying (1.3), or of semisimple local systems via harmonic bundles on the regular locus, without requiring any extra assumptions on the pluriharmonic metrics.
@@
Theorem 1.15 was proved for locally free $f$-numerically flat sheaves in the projective case [39, Theorem 1.2], building on the relative MMP for projective morphisms [4]. Its extension to the Kähler klt setting seems plausible, especially in light of recent work extending the relative MMP to analytic geometry [20, 29]. However, several technical results used in the proof of [38]—notably [49, Proposition 3.1], which appears in [38, Claim 4.5]—are, to the best of our knowledge, not directly available in the literature for the analytic setting. Moreover, in order to maintain our focus on Higgs sheaves and to keep the paper concise, we develop in Section 5.2 an alternative argument. It factors through local systems via an orbifold modification, uses the Jordan–Hölder filtration of the corresponding representation, and reconstructs the Higgs sheaf by descending the graded stable pieces based on Proposition 4.11.
Our overarching goal—proving Theorems 1.12 and 1.15—is the same as in [38, 39], the proofs we present in the Kähler setting are substantially different, for the reasons outlined above.
Acknowledgement. The authors wish to thank Xin Fu for useful discussions on uniform geometric estimates; Changpeng Pan and Di Wu for valuable conversations about harmonic bundles; and Masataka Iwai and Satoshi Jinnouchi for sharing their perspectives on quasi-uniformization problems. The second author acknowledges Tianshu Jiang for communications on related topics. The research was supported by the National Key R and D Program of China 2020YFA0713100. The authors are partially supported by NSF in China No.12141104, 12371062 and 12431004.
@@ 2.1. Complex orbifolds
A complex space $X$ with quotient singularities equipped with an effective orbifold structure $X_{\mathrm{orb}}:=\{V_{\alpha},G_{\alpha},\mu_{\alpha}\}$, namely, there exists an open cover $\{X_{\alpha}\}$ of $X$ such that $X_{\alpha}\cong V_{\alpha}/G_{\alpha}$ for some complex manifold $V_{\alpha}$ and $G_{\alpha}\subset\mathrm{Aut}(V_{\alpha})$ and $(V_{\alpha},G_{\alpha},\mu_{\alpha})$ are compatible along the overlaps, where $\mu_{\alpha}:V_{\alpha}\rightarrow X_{\alpha}$ is the quotient map induced by $G_{\alpha}$. Effective means that every $G_{\alpha}$-action is faithful. Indeed, $X$ always admits a standard orbifold structure, i.e. $\pi_{\alpha}$ is étale in codimension $1$. In virtue of puriry theorem of branch locus due to Grauert-Remmert [32], it can be easily seen that
Lemma 2.1 (c.f. [107, Proposition 1]).
Let $\{V_{\alpha},G_{\alpha},\pi_{\alpha}\}$ and $\{V_{\beta},G_{\beta},\pi_{\beta}\}$ be two standard orbifold structure. Then $V_{\alpha,\beta}:=\{(z_{\alpha},z_{\beta})\in V_{\alpha}\times V_{\beta}:\mu_{\alpha}(z_{\alpha})=\mu_{\beta}(z_{\beta})\}$ is smooth and the the natural projection $V_{\alpha,\beta}\rightarrow V_{\alpha}$ and $V_{\alpha,\beta}\rightarrow V_{\beta}$ are étale.
By definition, $V_{\alpha,\beta}/(G_{\alpha}\times G_{\beta})\cong X_{\alpha}\cap X_{\beta}$. Then by Zorn’s Lemma, there exists a standard orbifold structure that is maximal. When we consider a standard orbifold structure, we will always refer to the maximal one.
@@
Let $W_{orb}=(V_{i},G_{i})$ be a compact complex orbifold, and $\mathcal{E}_{orb}=\{\mathcal{E}_{i}\}$ be an torsion-free orbi-sheaf on $W_{orb}$. Then there exists an orbifold morphism $h_{orb}:Z_{orb}\rightarrow W_{orb}$ from a compact complex orbifold $Z_{orb}=\{(\widehat{V}_{i},G_{i},\widehat{\pi}_{i})\}$ so that the following holds.
@@
All objects (e.g. orbi sheaves, orbifold differential forms, etc.,) in the context of orbifolds are those defined on each $V_{\alpha}$ satisfying compatibility conditions. For a concise introduction to Higgs orbi-sheaves, we refer to [110, Section 2].
@@ 2.2. Higgs sheaves
We refer to [110, Section 4] for a concise discussion on Higgs sheaves over the regular locus of a compact complex normal space $X$. All discussions are still valid for locally free Higgs sheaves over $X$ by the following Remark 2.8.
@@ 2.2.1. Higgs sheaves over the whole space
We follow [37, Section 5] to introduce Higgs sheaves over normal spaces (see also [53, Section 3.2.2]).
@@
Let $\mathcal{E}_{X}$ be a locally free sheaf on a normal space $X$. Suppose that $\theta_{W}$ is a Higgs field of $\mathcal{E}_{X}|_{W}$ on an analytic subspace of codimension at least $2$, then $\theta_{W}$ extends to a Higgs field $\theta_{X}$ of $\mathcal{E}_{X}$.
@@
Recall that the saturation of $\theta_{X_{\mathrm{reg}}}$ torsion-free subsheaves increases the slope (see e.g. [110, Lemma 2.30]) and are again $\theta_{X_{\mathrm{reg}}}$-invariant by Lemma 2.5. Thus, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-stable if and only if
@@
holds for any saturated subsheaf $\mathcal{F}_{X}\subsetneq\mathcal{E}_{X}$ that is $\theta_{X_{\mathrm{reg}}}$-invariant. Ditto for semistability and polystability.
Combining Lemma 2.5 and Remark 2.6, it can be easily seen that
@@
Building on Lemma 2.4 and 2.7, all statements proved in the remainder of this subsection also hold for torsion-free Higgs sheaves over the whole space.
Following the spirits of [64, Pages 158], it can be easily seen that
@@
2.2.2. Pull-back of Higgs sheaves
The existence of the pull-back of Higgs sheaves is essential and relies on Kebekus-Schnell’s work [63] of constructing pull-back of reflexive differentials.
Lemma 2.10 (Reflexive pull-back of Higgs field, cf. [110, Lemma 4.9]).
Let $f:Y\rightarrow X$ be a holomorphism between complex spaces with klt singularities and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ be a reflexive Higgs sheaf on $X_{\mathrm{reg}}$. Let $\mathcal{E}_{Y_{\mathrm{reg}}}:=(f^{*}\mathcal{E}_{X})^{\vee\vee}|_{X_{\mathrm{reg}}}$ and $\sigma$ be the non-locally-fre locus of $\mathcal{E}_{X_{\mathrm{reg}}}$. If $f^{-1}(X_{\mathrm{reg}})\cap Y_{\mathrm{reg}}\neq\emptyset$, then $\mathcal{E}_{Y_{\mathrm{reg}}}$ admits a Higgs field $\theta_{Y_{\mathrm{reg}}}$ such that $\theta_{Y_{\mathrm{reg}}}=f^{*}\theta_{X_{\mathrm{reg}}}$ on $f^{-1}(X_{\mathrm{reg}}\setminus\Sigma)$.
Notation 2.11 (Reflexive pull-back of Higgs sheaves).
We call $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})$ the reflexive pull-back of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$, and denote it by $f^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$.
@@ 2.2.2. Pull-back of Higgs sheaves
If $\mathcal{E}_{X}$ is locally free, then $\mathcal{E}_{Y}=f^{*}\mathcal{E}_{X}$ is locally free and Lemma 2.4 implies $\theta_{Y_{\mathrm{reg}}}$ extends to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}$. We call $(\mathcal{E}_{Y},\theta_{Y})$ the pull-back of $(\mathcal{E}_{X},\theta_{X})$, and denote it by $f^{*}(\mathcal{E}_{X},\theta_{X})$.
@@ 2.3. Orbifold Chern classes
We refer to [90, Section 2] for a concise introduction of the definition of orbifold Chern classes via partial modifications [67]. A compact complex klt space has only quotient singularities in codimension 2 (cf. [31, Lemma 5.8]) and so admits a partial orbifold modification by [67], which allows us to define the orbifold first and second Chern classes as below.
Definition 2.13 (cf. [90, Section 2]).
Let $X$ be a compact complex klt space, $f:Y\rightarrow X$ be a partial orbifold resolution and $Y_{\mathrm{orb}}=\{V_{i},G_{i},\pi_{i}\}$ be the standard orbifold structure of $Y$. For any reflexive sheaf $\mathcal{E}$, $\widehat{c}_{2}(\mathcal{E})\in H_{2n-4}(X,\mathbb{R})$ is defined by
$\widehat{c}_{2}(\mathcal{E})\cdot\sigma:=c_{2}^{\mathrm{orb}}(f^{[*]}_{\mathrm{orb}}\mathcal{E})\cdot f^{*}\sigma,\ \forall\sigma\in H^{2n-4}(X,\mathbb{R})$
where $f^{[*]}_{\mathrm{orb}}\mathcal{E}=\{(f\circ\pi_{i})^{[*]}\mathcal{E}\}$ is a reflexive orbi-sheaf on $X_{\mathrm{orb}}$. Analogously, $\widehat{c}_{1}(\mathcal{E})\in{H_{2n-2}(X,\mathbb{R})}$ and $\widehat{c}_{1}^{2}(\mathcal{E})\in H_{2n-4}(X,\mathbb{R})$ can be introduced.
Definition 2.13 is independent of the choice of $f$. we refer to [90, Section 2] for a detailed discussion. When $\mathcal{E}=\mathcal{T}_{X}$, it coincides with the definition of $\widehat{c}_{2}(X)$ in [75, 31, 14].
@@
Here, $c_{2}^{\mathrm{orb}}(\cdot)$ denotes the orbifold Chern character in Bott–Chern cohomology for general coherent orbifold sheaves, introduced in [78]; for vector bundles this coincides with the definition via metrics given in [77]. The Grothendieck–Riemann–Roch formula for orbifold embeddings is also established in [78].
The following proposition is analogous to [55, Lemma 2.13] for $\mathbb{Q}$-Chern classes.
@@
and if $\omega_{0},\cdots,\omega_{n-2}$ are Kähler and the equality in (2.2) holds then the natural map $\mathcal{Q}\rightarrow\mathcal{Q}^{\vee\vee}$ is an isomorphism in codimension $2$.
Proof of Proposition 2.15 essentially follows by considering the equality case of the argument of [53, Proposition 4.23] concerning $\widehat{c}_{2}(\mathcal{E})$ (cf. [62, Lemma 3.2] and [55, Lemma 2.13] for the projective case). We recall the following necessary lemmas in our proof.
Lemma 2.16 (cf. [90, Lemma 2.2]).
Let $f:Y\rightarrow X$ be a projective bimeromorphism between compact complex spaces of dimension $n$. Suppose that $\sigma_{1},\sigma_{2}\in H_{2n-2k}(X,\mathbb{R})$ agrees outside $f^{-1}(Z)$ for some analytic subspace $Z\subset X$ of codimension at least $k$. Then $f_{*}\sigma_{1}=f_{*}\sigma_{2}\in H_{2n-2k}(Y,\mathbb{R})$.
Lemma 2.17 (cf. [110, Lemma 2.17]).
Let $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow X_{\mathrm{orb}}$ be a bimeromorphism between compact complex orbifolds. Suppose that $\mathcal{F}_{\mathrm{orb}},\mathcal{E}_{\mathrm{orb}}$ are coherent orbi-sheaves on $Y_{\mathrm{orb}},X_{\mathrm{orb}}$, respectively, such that $(g_{\mathrm{orb}})_{*}\mathcal{F}_{\mathrm{orb}}=\mathcal{E}_{\mathrm{orb}}$ in codimension $k$ and the indeterminacy locus of $g_{\mathrm{orb}}^{-1}$ has codimension at least $k+1$. Then $g_{*}\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}})=\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}}),$ i.e. $\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}})\cdot g^{*}\gamma=\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}})\cdot\gamma$ for any $\gamma\in H^{2n-2k}(X,\mathbb{R})$,
@@
Applying Lemma 2.2 and 2.17 and adapt the notations of Lemma 2.2 and Definition 2.13, it holds that
$\forall\sigma\in H^{2n-4}(X,\mathbb{R}),\ \widehat{c}_{2}(\mathcal{E})\cdot\sigma:=c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})\cdot g^{*}f^{*}\sigma.$
Ditto for $\widehat{c}_{1}(\mathcal{E})$ and $\widehat{c}_{1}^{2}(\mathcal{E})$.
@@
Suppose that $X$ is a compact complex normal space of dimension $n$ equipped with a Kähler form $\omega_{X}$ and $f:Y\rightarrow X$ is a bimeromorphism from a compact complex space $Y$ with only quotient singularities to $X$. Let $Z$ be the indeterminacy locus of $f^{-1}$ and $Y_{\mathrm{orb}}:=\{(V_{i},G_{i},\mu_{i})\}$ be a complex orbifold structure. Then for any $V\subset\subset X\setminus Z$, there exists an smooth orbifold semi-positive $(1,1)$-form $\omega_{V,\mathrm{orb}}\in[f^{*}_{\mathrm{orb}}\omega_{X}]\in H^{1,1}_{BC}(X_{\mathrm{orb}},\mathbb{R})$ such that $\omega_{V,\mathrm{orb}}$ is strictly positive on $f_{\mathrm{orb}}^{*}V:=\{(f\circ\pi_{i})^{*}U\}$.
This lemma follows from [7, Section 6.1]. Note that let $f$ be the identity map $Y\rightarrow Y$, we conclude that $Y_{\mathrm{orb}}$ is Kähler if $Y$ is Kähler.
@@
Let $g:W\rightarrow X$ be a partial orbifold resolution by Lemma 2.2 and $W_{\mathrm{orb}}:=\{(V_{i},G_{i},\pi_{i})\}$ be the standard orbifold structure of $W$. Denote the indeterminacy locus of $g^{-1}$ by $Z$. Let $g_{i}:=g\circ\pi_{i}$, $\mathcal{E}_{\mathrm{orb}}:=g_{\mathrm{orb}}^{[*]}\mathcal{E}$ and $\mathcal{F}_{\mathrm{orb}}$ be the saturation of $\{\operatorname{image}(g_{i}^{[*]}\mathcal{F}\rightarrow g_{i}^{[*]}\mathcal{E})\}$, where $g_{i}^{[*]}\mathcal{F}\rightarrow g_{i}^{[*]}\mathcal{E}$ is induced by the inclusion morphism $\mathcal{F}\rightarrow\mathcal{E}$. Since $\pi_{i}$ is étale in codimension $1$, $\mathcal{F}_{\mathrm{orb}}=g_{\mathrm{orb}}^{[*]}\mathcal{F}$ outside $g_{\mathrm{orb}}^{-1}(Z)$ due to the reflexivity. Then we have

### v2

@@ 1.3. Overview and main technical components
The careful treatment of harmonic bundles on the regular locus presented above is necessary in the Kähler klt context because Mochizuki’s theory of tame harmonic bundles on quasi-projective manifolds [84, 85, 86]—which crucially underlies the construction of $\mu_{X_{\mathrm{reg}}}$ in the projective case [39]—has not yet been established for general quasi-Kähler manifolds.
It should be noted that the admissibility of the limiting Hermitian-Einstein metric constructed in Section 3 remains unknown (cf. Question 3.6). The key point of Steps 2 and 3 is that klt singularities are sufficiently mild to allow us, via harmonic bundles on the regular locus, to characterize polystable Higgs bundles satisfying (1.3) and semisimple local systems on the regular locus, without imposing any extra assumptions on the pluriharmonic metrics.
@@
Theorem 1.15 was proved for locally free $f$-numerically flat sheaves in the projective case [39, Theorem 1.2], building on the relative MMP for projective morphisms [4]. Its extension to the Kähler klt setting appears plausible, especially in light of recent work extending the relative MMP to the analytic context [20, 29]. However, several technical results used in the proof of [38]—notably [50, Proposition 3.1], which appears in the proof of [38, Claim 4.5]—do not seem to be directly available in the analytic setting, to the best of our knowledge and based on the existing literature. Moreover, in order to maintain our focus on Higgs sheaves and to keep the paper concise, we adopt an alternative argument in Section 5.2. This argument proceeds via an orbifold modification to factor through local systems, employs Jordan-Hölder filtrations of the corresponding representation, and reconstructs the Higgs sheaf by descending the graded stable pieces based on Proposition 4.11.
Although our overarching goal—proving Theorems 1.12 and 1.15—coincides with that of [38], the proofs we present in the Kähler setting are substantially different for the reasons outlined above.
Acknowledgement. The authors thank Xin Fu for insightful discussions on uniform geometric estimates, Osamu Fujino for kind help and sharing the preprint [30] with them, Changpeng Pan and Di Wu for helpful conversations regarding harmonic bundles, and Masataka Iwai and Satoshi Jinnouchi for valuable remarks on quasi-uniformization problems. The second author also thanks Tianshu Jiang and Shin-ichi Matsumura for conversations in related topics. The research was supported by the National Key R and D Program of China 2020YFA0713100. The authors are partially supported by NSF in China No.12141104, 12371062 and 12431004.
@@ 2.1. Complex orbifolds
A complex space $X$ with quotient singularities equipped with an effective orbifold structure $X_{\mathrm{orb}}:=\{(V_{\alpha},G_{\alpha},\mu_{\alpha})\}$, namely, there exists an open cover $\{X_{\alpha}\}$ of $X$ such that $X_{\alpha}\cong V_{\alpha}/G_{\alpha}$ for some complex manifold $V_{\alpha}$ and $G_{\alpha}\subset\mathrm{Aut}(V_{\alpha})$ and $(V_{\alpha},G_{\alpha},\mu_{\alpha})$ are compatible along the overlaps, where $\mu_{\alpha}:V_{\alpha}\rightarrow X_{\alpha}$ is the quotient map induced by $G_{\alpha}$. Effective means that every $G_{\alpha}$-action is faithful. Indeed, $X$ always admits a standard orbifold structure, i.e., $\pi_{\alpha}$ is étale in codimension $1$. In virtue of purity theorem of branch locus due to Grauert-Remmert [32], it can be easily seen that
Lemma 2.1 (c.f. [110, Proposition 1]).
Let $\{(V_{\alpha},G_{\alpha},\mu_{\alpha})\}$ and $\{(V_{\beta},G_{\beta},\mu_{\beta})\}$ be two standard orbifold structures. Then $V_{\alpha,\beta}:=\{(z_{\alpha},z_{\beta})\in V_{\alpha}\times V_{\beta}:\mu_{\alpha}(z_{\alpha})=\mu_{\beta}(z_{\beta})\}$ is smooth and the natural projections $V_{\alpha,\beta}\rightarrow V_{\alpha}$ and $V_{\alpha,\beta}\rightarrow V_{\beta}$ are étale.
The definition gives $V_{\alpha,\beta}/(G_{\alpha}\times G_{\beta})\cong X_{\alpha}\cap X_{\beta}$. Then according to Zorn’s Lemma, there exists a standard orbifold structure that is maximal. By a standard orbifold structure we will always mean the maximal one.
@@
Let $W_{orb}=\{(V_{i},G_{i},\mu_{i})\}$ be a compact complex orbifold, and $\mathcal{E}_{orb}=\{\mathcal{E}_{i}\}$ be a torsion-free orbi-sheaf on $W_{orb}$. Then there exists an orbifold morphism $h_{orb}:Z_{orb}\rightarrow W_{orb}$ from a compact complex orbifold $Z_{orb}=\{(\widehat{V}_{i},G_{i},\widehat{\mu}_{i})\}$ so that the following holds:
@@
All objects (e.g., orbi sheaves, orbifold differential forms, etc.) in the context of orbifolds are those defined on each $V_{\alpha}$ that satisfying compatibility conditions. For a concise introduction to Higgs orbi-sheaves, we refer to [113, Section 2].
@@ 2.2. Higgs sheaves
We refer to [113, Section 4] for a concise discussion on Higgs sheaves over the regular locus of a compact complex normal space $X$. All discussions are still valid for locally free Higgs sheaves over $X$ by the following Remark 2.8.
@@ 2.2.1. Higgs sheaves over the whole space
We follow [37, Section 5] to introduce Higgs sheaves over normal spaces (see also [54, Section 3.2.2]).
@@
Let $\mathcal{E}_{X}$ be a locally free sheaf on a normal space $X$. Suppose that $\theta_{W}$ is a Higgs field of $(\mathcal{E}_{X})|_{W}$ on an analytic subspace $W$ of codimension at least $2$, then $\theta_{W}$ extends to a Higgs field $\theta_{X}$ of $\mathcal{E}_{X}$.
@@
Recall that the saturation of a $\theta_{X_{\mathrm{reg}}}$-invariant torsion-free subsheaf increases the slope (see e.g. [113, Lemma 2.30]) and is again $\theta_{X_{\mathrm{reg}}}$-invariant by Lemma 2.5. Thus, for any nef polarization $(\alpha_{0},\cdots,\alpha_{n-2})$, $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\alpha_{0},\cdots,\alpha_{n-2})$-stable if and only if
@@
holds for any saturated subsheaf $\mathcal{F}_{X}\subsetneq\mathcal{E}_{X}$ that is $\theta_{X_{\mathrm{reg}}}$-invariant. Smilar statements hold for semistability and polystability.
Combining Lemma 2.5 and Remark 2.6 shows that
@@
Building on Lemmas 2.4 and 2.7, all statements proved in the remainder of this subsection also hold for torsion-free Higgs sheaves over the whole space.
Following the spirit of [65, Page 158], it can be easily seen that
@@
2.2.2. Pullback of Higgs sheaves
The existence of the pullback of Higgs sheaves is essential and relies on Kebekus-Schnell’s work [64] on constructing the pullback of reflexive differentials.
Lemma 2.10 (Reflexive pullback of Higgs field, cf. [113, Lemma 4.9]).
Let $f:Y\rightarrow X$ be a holomorphism between complex spaces with klt singularities and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ be a reflexive Higgs sheaf on $X_{\mathrm{reg}}$. Let $\mathcal{E}_{Y_{\mathrm{reg}}}:=(f^{*}\mathcal{E}_{X})^{\vee\vee}|_{Y_{\mathrm{reg}}}$ and $\Sigma$ be the non-locally-free locus of $\mathcal{E}_{X_{\mathrm{reg}}}$. If $f^{-1}(X_{\mathrm{reg}})\cap Y_{\mathrm{reg}}\neq\emptyset$, then $\mathcal{E}_{Y_{\mathrm{reg}}}$ admits a Higgs field $\theta_{Y_{\mathrm{reg}}}$ such that $\theta_{Y_{\mathrm{reg}}}=f^{*}\theta_{X_{\mathrm{reg}}}$ on $f^{-1}(X_{\mathrm{reg}}\setminus\Sigma)$.
Notation 2.11 (Reflexive pullback of Higgs sheaves).
We call $(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})$ the reflexive pullback of $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$, and denote it by $f^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$.
@@ 2.2.2. Pullback of Higgs sheaves
If $\mathcal{E}_{X}$ is locally free, then $\mathcal{E}_{Y}=f^{*}\mathcal{E}_{X}$ is locally free and Lemma 2.4 implies $\theta_{Y_{\mathrm{reg}}}$ extends to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}$. We call $(\mathcal{E}_{Y},\theta_{Y})$ the pullback of $(\mathcal{E}_{X},\theta_{X})$, and denote it by $f^{*}(\mathcal{E}_{X},\theta_{X})$.
@@ 2.3. Orbifold Chern classes
We refer to [92, Section 2] for a concise introduction to the definition of orbifold Chern classes via orbifold modifications [68]. A compact complex klt space $X$ has only quotient singularities in codimension $2$ (cf. [31, Lemma 5.8]) and therefore admits an orbifold modification by [68], i.e., there exists a bimeromorphic map $f:Y\rightarrow X$ from a complex space with only quotient singularities to $X$ such that the indeterminacy locus of $f^{-1}$ has codimension at least $3$. This allows us to define the orbifold first and second Chern classes as follows.
Definition 2.13 (cf. [92, Section 2]).
Let $X$ be a compact complex klt space, $f:Y\rightarrow X$ be an orbifold modification and $Y_{\mathrm{orb}}=\{(V_{i},G_{i},\mu_{i})\}$ be the standard orbifold structure of $Y$. For any reflexive sheaf $\mathcal{E}$ on $X$, $\widehat{c}_{2}(\mathcal{E})\in H_{2n-4}(X,\mathbb{R})$ is defined by
$\widehat{c}_{2}(\mathcal{E})\cdot\sigma:=c_{2}^{\mathrm{orb}}(f^{[*]}_{\mathrm{orb}}\mathcal{E})\cdot f^{*}\sigma,\ \forall\ \sigma\in H^{2n-4}(X,\mathbb{R}),$
where $f^{[*]}_{\mathrm{orb}}\mathcal{E}=\{(f\circ\mu_{i})^{[*]}\mathcal{E}\}$ is a reflexive orbi-sheaf on $Y_{\mathrm{orb}}$. The classes $\widehat{c}_{1}(\mathcal{E})\in{H_{2n-2}(X,\mathbb{R})}$ and $\widehat{c}_{1}^{2}(\mathcal{E})\in H_{2n-4}(X,\mathbb{R})$ are defined analogously.
Definition 2.13 is independent of the choice of $f$. We refer to [92, Section 2] for a detailed discussion. When $\mathcal{E}=\mathcal{T}_{X}$, it coincides with the definition of $\widehat{c}_{2}(X)$ in [76, 31, 14] (see also [42]).
@@
Here, $c_{2}^{\mathrm{orb}}(\cdot)$ denotes the orbifold Chern class in Bott-Chern cohomology for general coherent orbifold sheaves, introduced in [79]; for vector bundles this coincides with the definition via metrics given in [78]. The Grothendieck-Riemann-Roch formula for orbifold embeddings is established in [79].
The following proposition is analogous to [56, Lemma 2.13] for $\mathbb{Q}$-Chern classes.
@@
where $\widehat{ch}_{2}(\cdot):=\frac{1}{2}\widehat{c}_{1}^{2}(\cdot)-\widehat{c}_{2}(\cdot).$ If $\omega_{0},\cdots,\omega_{n-2}$ are Kähler forms and the equality in (2.2) holds, then the natural map $\mathcal{Q}\rightarrow\mathcal{Q}^{\vee\vee}$ is an isomorphism in codimension $2$.
Proof of Proposition 2.15 essentially follows by considering the equality case of the argument of [54, Proposition 4.23] concerning $\widehat{c}_{2}(\mathcal{E})$ (see [63, Lemma 3.2] and [56, Lemma 2.13] for the projective case). We recall the following necessary lemmas in our proof.
Lemma 2.16 (cf. [92, Lemma 2.2]).
Let $f:Y\rightarrow X$ be a projective bimeromorphism between compact complex spaces of dimension $n$. Suppose that $\sigma_{1},\sigma_{2}\in H_{2n-2k}(Y,\mathbb{R})$ agrees outside $f^{-1}(Z)$ for some analytic subspace $Z\subset X$ of codimension at least $k$. Then $f_{*}\sigma_{1}=f_{*}\sigma_{2}\in H_{2n-2k}(X,\mathbb{R})$.
Lemma 2.17 (cf. [113, Lemma 2.17]).
Let $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow X_{\mathrm{orb}}$ be an orbifold bimeromorphism between compact complex orbifolds. Suppose that $\mathcal{F}_{\mathrm{orb}}$ and $\mathcal{E}_{\mathrm{orb}}$ are coherent orbi-sheaves on $Y_{\mathrm{orb}}$ and $X_{\mathrm{orb}}$, respectively, such that $(g_{\mathrm{orb}})_{*}\mathcal{F}_{\mathrm{orb}}=\mathcal{E}_{\mathrm{orb}}$ in codimension $k$ and the indeterminacy locus of $g_{\mathrm{orb}}^{-1}$ has codimension at least $k+1$. Then $g_{*}\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}})=\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}})$; that is, $\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{F}_{\mathrm{orb}})\cdot g^{*}\gamma=\mathrm{ch}_{k}^{\mathrm{orb}}(\mathcal{E}_{\mathrm{orb}})\cdot\gamma$ for any $\gamma\in H^{2n-2k}(X,\mathbb{R})$, where $g$ is the induced morphism between the underlying complex spaces $Y$ and $X$ of $Y_{\mathrm{orb}}$ and $X_{\mathrm{orb}}$, respectively.
@@
Let the notation be as in Definition 2.13. For a reflexive sheaf $\mathcal{E}$ on $X$, apply Lemma 2.2 to the reflexive orbi-sheaf $\mathcal{E}_{\mathrm{orb}}:=f^{[*]}_{\mathrm{orb}}\mathcal{E}$ on $Y_{\mathrm{orb}}$ to obtain an orbifold morphism $h_{orb}:Z_{orb}\rightarrow Y_{orb}$ such that $E_{\mathrm{orb}}=h_{\mathrm{orb}}^{*}\mathcal{E}_{\mathrm{orb}}/\mathrm{tor}:=\{h_{i}^{*}\mathcal{E}_{i}/\mathrm{tor}\}$ is a vector orbi-bundle on $Z_{\mathrm{orb}}$. Then for any $\sigma\in H^{2n-4}(X,\mathbb{R})$,
[…]
where $h:Z\rightarrow Y$ is the induced map on underlying complex spaces. The same holds for $\widehat{c}_{1}(\mathcal{E})$ and $\widehat{c}_{1}^{2}(\mathcal{E})$.
@@
Suppose that $X$ is a compact complex normal space of dimension $n$ equipped with a Kähler form $\omega_{X}$ and $f:Y\rightarrow X$ is a bimeromorphism from a compact complex space $Y$ with only quotient singularities to $X$. Let $Z$ be the indeterminacy locus of $f^{-1}$ and $Y_{\mathrm{orb}}:=\{(V_{i},G_{i},\mu_{i})\}$ be a complex orbifold structure of $Y$. Then for any $V\subset\subset X\setminus Z$, there exists a smooth orbifold semipositive $(1,1)$-form $\omega_{V,\mathrm{orb}}\in[f^{*}_{\mathrm{orb}}\omega_{X}]\in H^{1,1}_{\rm{BC}}(Y_{\mathrm{orb}},\mathbb{R})$ such that $\omega_{V,\mathrm{orb}}$ is strictly positive on $f_{\mathrm{orb}}^{-1}V:=\{(f\circ\mu_{i})^{-1}V\}$.
This lemma follows from [7, Section 6.1]. In particular, taking $f$ to be the identity map $Y\rightarrow Y$, we conclude that $Y_{\mathrm{orb}}$ is Kähler if $Y$ is Kähler.
@@
Let $g:W\rightarrow X$ be a partial orbifold resolution by Lemma 2.2 and $W_{\mathrm{orb}}:=\{(V_{i},G_{i},\mu_{i})\}$ be the standard orbifold structure of $W$. Denote by $Z$ the indeterminacy locus of $g^{-1}$. Let $g_{i}:=g\circ\mu_{i}$, $\mathcal{E}_{\mathrm{orb}}:=g_{\mathrm{orb}}^{[*]}\mathcal{E}$ and $\mathcal{F}_{\mathrm{orb}}$ be the saturation of $\{\operatorname{image}(g_{i}^{[*]}\mathcal{F}\rightarrow g_{i}^{[*]}\mathcal{E})\}$, where $g_{i}^{[*]}\mathcal{F}\rightarrow g_{i}^{[*]}\mathcal{E}$ is induced by the inclusion morphism $\mathcal{F}\rightarrow\mathcal{E}$. Since $\mu_{i}$ is étale in codimension $1$, $\mathcal{F}_{\mathrm{orb}}=g_{\mathrm{orb}}^{[*]}\mathcal{F}$ outside $g_{\mathrm{orb}}^{-1}(Z)$ due to reflexivity. Then we have